OpenStudy (anonymous):

Which of the following is the slope between the points (-6, 0) and (2, -9)? 8/9, -8/9, 9/8, or -9/8

OpenStudy (anonymous):

Slope of a line passing through two points (x1,y1) and (x2,y2) = (y1-y2)/(x1-x2) Can you use this info ?

OpenStudy (kirbykirby):

For 2 points \( (x_1, y_1)\) and \((x_2, y_2)\), the slope is: \[ \text{slope} =\frac{\Delta y}{\Delta x}=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

i did not get any of the answers

OpenStudy (kirbykirby):

say \((-6, 0)\) is \((x_1, y_1)\) and \((2, -9)\) is \((x_2, y_2)\)

OpenStudy (anonymous):

i get 4/9

OpenStudy (anonymous):

You should get -9/8